Tom is 10 years older than Pat. In 6 years. Tom will be 1.5 times as old as Pat will be then. Find their present ages.
Answer by oberobic(2304)

T = P + 10
.
In 6 years,
T+6
P+6
.
T+6 = 1.5(P+6)
.
Substitute for T
.
(P+10) + 6 = 1.5P + 9
P +16 = 1.5P + 9
-.5P = -7
P = 14 = Pat's age
.
T = P+10 = Tom's age
T = 24
.
Check the values in 6 years.
.
30 = 1.5(20) ?
Yes
.
Answer: Pat is 14 and Tom is 24.
.
Done.